OpenStudy (yanasidlinskiy):

Relations and Functions y=-4x+3 Complete the table of values for the equation: x: -3, -1, 0, 4 y:

OpenStudy (yanasidlinskiy):

My first step would be: -4(-3) + 3 12+3 Right?

ganeshie8 (ganeshie8):

Yes!

OpenStudy (yanasidlinskiy):

Then... -4(-1) + 3 4+3 7 Right?

ganeshie8 (ganeshie8):

Looks good! I'm sure you're going to complete rest of the table just fine

OpenStudy (yanasidlinskiy):

-4(0) + 3 0+3 3 -4(4) + 3 -16+3 -13

OpenStudy (yanasidlinskiy):

Hopefully, ganeshie! I hope it's correct lol.

ganeshie8 (ganeshie8):

All correct. good job !

ganeshie8 (ganeshie8):

As you can see we just need to follow the simple rules

OpenStudy (yanasidlinskiy):

Wait, wait! I think I'm putting the equation up wrong.... Like on my actual worksheet it's y-4x=3 This equation would change all the answers, right? If it will then I did it wrong. Idk why I was switching up everything.

ganeshie8 (ganeshie8):

Yes scratch everything and start over

ganeshie8 (ganeshie8):

\[y-4x~=~3\] is same as \[y~=~4x+3\]

OpenStudy (yanasidlinskiy):

How?

OpenStudy (yanasidlinskiy):

The 4x is negative... How would it be positive?

ganeshie8 (ganeshie8):

We're given below equation\[y-4x~=~3\] Since it is an equation, adding something to "both" sides won't change it, yes ?

ganeshie8 (ganeshie8):

add \(4x\) both sides : \[y-4x\color{red}{+4x}~=~3\color{red}{+4x}\]

OpenStudy (yanasidlinskiy):

Ohhhh, ok! I see! So like even if I had an equation of lets say: y-8x +5 I would get: y= 8x+5 Right?

ganeshie8 (ganeshie8):

thats right!

OpenStudy (yanasidlinskiy):

Ahh, ok! What about like if I had lets say umm... y+10x - 5 I'd get: y = -10x - 5 Right?

ganeshie8 (ganeshie8):

Is it \(y+10x-5\) or \(y+10x=5\) ?

OpenStudy (yanasidlinskiy):

y+10x= -5 My bad for the confusion.

OpenStudy (yanasidlinskiy):

It'd be y = -10x - 5. Is that correct?

ganeshie8 (ganeshie8):

Correct! Notice that you're simply subtracting \(10x\) both sides to get that

ganeshie8 (ganeshie8):

y+10x= -5 subtract 10x both sides y+10x-10x = -5-10x y = -5-10x

OpenStudy (yanasidlinskiy):

Yes, yes! Wow! I honestly don't even know what I would have done without you! Wow!! I'm soooo happy I understand this now!!

ganeshie8 (ganeshie8):

takes time, follow the rules and you will be fine :)

OpenStudy (yanasidlinskiy):

Yes, definitely. Like I knew I had to do something with the equation, I just didn't know why and how. Like I understand this so much better!!! Thank you!!!

ganeshie8 (ganeshie8):

the most important rule with equations is that you're allowed to add/subtract "same thing" to both sides of an equation.

ganeshie8 (ganeshie8):

np :)

OpenStudy (yanasidlinskiy):

Yes, definitely!

OpenStudy (yanasidlinskiy):

Well, guess who I'll be coming to next time I need help;) You know it;)

ganeshie8 (ganeshie8):

sure... any time !

OpenStudy (yanasidlinskiy):

Alrighty! Thanks again! I appreciate it so much! As of now, I do have to go and move on to the next level of topic.

ganeshie8 (ganeshie8):

just an example, if ``` yana = banana ``` then ``` yana + apple = banana + apple ``` you're allowed to add same thing (apple) to both sides. above both are called equations because they have a "=" symbol in them. both the equations are same.

OpenStudy (yanasidlinskiy):

LOL, haha! I'm so going to save that haha! Thanks for that example;)